# Adversarial Training Should Be Cast as a Non-Zero-Sum Game

### v1

@@ Abstract
One prominent approach toward resolving the adversarial vulnerability of deep neural networks is the two-player zero-sum paradigm of adversarial training, in which predictors are trained against adversarially-chosen perturbations of data. Despite the promise of this approach, algorithms based on this paradigm have not engendered sufficient levels of robustness, and suffer from pathological behavior like robust overfitting. To understand this shortcoming, we first show that the commonly used surrogate-based relaxation used in adversarial training algorithms voids all guarantees on the robustness of trained classifiers. The identification of this pitfall informs a novel non-zero-sum bilevel formulation of adversarial training, wherein each player optimizes a different objective function. Our formulation naturally yields a simple algorithmic framework that matches and in some cases outperforms state-of-the-art attacks, attains comparable levels of robustness to standard adversarial training algorithms, and does not suffer from robust overfitting.
@@ 1 Introduction
A longstanding disappointment in the machine learning (ML) community is that deep neural networks (DNNs) remain vulnerable to seemingly innocuous changes to their input data including nuisances in visual data Robey et al. (2020); Laidlaw et al. (2020); Hendrycks and Dietterich (2019); Hendrycks et al. (2021); Eykholt et al. (2018), sub-populations Santurkar et al. (2021); Sohoni et al. (2020); Koh et al. (2021); Zhou et al. (2022), and distribution shifts Xiao et al. (2021); Arjovsky et al. (2019); Eastwood et al. (2022); Sagawa et al. (2020); Robey et al. (2021a). Prominent amongst these vulnerabilities is the setting of adversarial examples, wherein it has been conclusively shown that imperceptible, adversarially-chosen perturbations can fool state-of-the-art classifiers parameterized by DNNs (Szegedy et al., 2013; Biggio et al., 2013a, 2012; Carlini and Wagner, 2017; Biggio et al., 2013b). In response, a plethora of research has proposed so-called adversarial training (AT) algorithms (Huang et al., 2015; Wong and Kolter, 2018; Kurakin et al., 2017; Robey et al., 2021b; Sridhar et al., 2022; Madry et al., 2018; Goodfellow et al., 2015), which are designed to improve robustness against adversarial examples.
@@
AT is ubiquitously formulated as a two-player zero-sum game, where both players—often referred to as the defender and the adversary—respectively seek to minimize and maximize the classification error. However, this zero-sum game is not implementable in practice as the discontinuous nature of the classification error is not compatible with first-order optimization algorithms. To bridge this gap between theory and practice, it is commonplace to replace the classification error with a smooth surrogate loss (e.g., the cross-entropy loss) which is amenable to gradient-based optimization Madry et al. (2018); Zhang et al. (2019). And while this seemingly harmless modification has a decades-long tradition in the ML literature due to the guarantees it imparts on non-adversarial objectives Bartlett et al. (2006); Shalev-Shwartz and Ben-David (2014); Roux (2017), there is a pronounced gap in the literature regarding the implications of this relaxation on the standard formulation of AT.
@@
As the field of robust ML has matured, surrogate-based AT algorithms (see, e.g., Madry et al. (2018); Zhang et al. (2019); Goodfellow et al. (2015); Wang et al. (2020)) have collectively ushered in significant progress toward designing stronger attacks and obtaining more robust defenses Croce et al. (2020a). However, despite these advances, recent years have witnessed a plateau in robustness measures on leaderboards such as RobustBench, resulting in the widely held beliefs that robustness and accuracy may be irreconcilable Tsipras et al. (2019a); Dobriban et al. (2020); Javanmard et al. (2020) and that robust generalization requires significantly more data Schmidt et al. (2018); Chen et al. (2020); Stutz et al. (2019). Moreover, various phenomena such as robust overfitting Rice et al. (2020) and insufficient robustness evaluation Carlini et al. (2019) have indicated that progress has been overestimated Croce and Hein (2020). To combat these pitfalls, state-of-the-art algorithms increasingly rely on ad-hoc regularization schemes Kannan et al. (2018); Zhang et al. (2019); Chan et al. (2020); Hoffman et al. (2019); Finlay et al. (2018), weight perturbations Wu et al. (2020); Sun et al. (2021); Foret et al. (2020), and heuristics such as multiple restarts Madry et al. (2018), carefully crafted learning rate schedules Rice et al. (2020), and convoluted stopping conditions Croce and Hein (2020), all of which contribute to an unclear set of best practices and a growing literature concerned with identifying flaws in various AT schemes Latorre et al. (2023).
@@
Motivated by these challenges, we argue that the pervasive surrogate-based zero-sum approach to AT suffers from a fundamental flaw. Our analysis of the standard minimax formulation of AT reveals that maximizing a surrogate like the cross-entropy provides no guarantee that the the classification error will increase, resulting in weak adversaries and ineffective AT algorithms. In identifying this shortcoming, we prove that to preserve guarantees on the optimality of the classification error objective, the defender and the adversary must optimize different objectives, resulting in a non-zero-sum game. This leads to a novel, yet natural bilevel formulation (Bard, 2013) of AT in which the defender minimizes an upper bound on the classification error, while the attacker maximizes a continuous reformulation of the classification error. We then propose an algorithm based on our formulation which is free from heuristics and ad hoc optimization techniques. Our empirical evaluations reveal that our approach matches the test robustness achieved by the state-of-the-art, yet highly heuristic approaches such as AutoAttack, and that it eliminates the problem of robust overfitting.
Contributions. We summarize our contributions in the following bullets.
@@
  • New adversarial training algorithm. We derive a new, heuristic-free algorithm (Algorithm 2) based on our bilevel formulation, and show that offers strong robustness on CIFAR-10.
  • Elimination of robust overfitting. Without the need of heuristic modifications, our algorithm does not suffer from robust overfitting (RO). This suggest that RO is an artifact of the use of improper surrogates in the original AT paradigm, and that the use of a correct optimization formulation is enough to solve it.
  • State-of-the-art robustness evaluation. We show that our proposed optimization objective for the adversary yields a simple algorithm that matches the performance of the state-of-the-art, yet highly complex AutoAttack method, on classifiers trained on CIFAR-10.
@@ 2.1 Preliminaries: Training DNNs with surrogate losses
Prominent among the barriers to solving (1) in practice is the fact that the classification error is a discontinuous function of $\theta$, which in turn renders continuous first-order methods (e.g., gradient descent) intractable. Fortunately, this pitfall can be resolved by minimizing a surrogate loss function $\ell:[k]\times[k]\to\mathbb{R}$ in place of the classification error (Shalev-Shwartz and Ben-David, 2014, §12.3). For minimization problems, surrogate losses are chosen to be differentiable upper bounds of the classification error of $f_{\theta}$, in the sense that
@@ 2.2 The pervasive setting of adversarial examples
For common hypothesis classes, it is well-known that classifiers obtained by solving (3) are sensitive to seemingly benign changes to their input data. Among these vulnerabilities, perhaps the most well-studied is the setting of adversarial examples, wherein a plethora of research has demonstrated that state-of-the-art classifiers can be fooled by small, adversarially-chosen perturbations (Szegedy et al., 2013; Biggio et al., 2013a, 2012; Carlini and Wagner, 2017; Biggio et al., 2013b). In other words, given an instance label pair $(X,Y)$, it is relatively straightforward to find perturbations $\eta\in\mathbb{R}^{d}$ with small norm $\left|\left|\eta\right|\right|\leq\epsilon$ for some fixed $\epsilon>0$ such that
@@ 2.3 Surrogate-based approaches to robustness
As discussed in § 2.1, the discontinuity of the classification error complicates the task of finding adversarial examples, as in (5), and of training against these perturbed instances, as in (6). One appealing approach toward overcoming this pitfall is to simply deploy a surrogate loss in place of the classification error inside (6), which gives rise to the following pair of optimization problems:
$\eta^{\star}\in\operatorname*{arg\,max}_{\eta:\left|\left|\eta\right|\right|\leq\epsilon}\ell(f_{\theta}(X+\eta),Y)$ (7)
$\min_{\theta\in\Theta}\mathbb{E}\left[\max_{\eta:\|\eta\|\leq\epsilon}\ell(f_{\theta}(X+\eta),Y)\right]$ (8)
Indeed, this surrogate-based approach is pervasive in practice. Madry et al.’s seminal paper on the subject of adversarial training employs this formulation (Madry et al., 2018), which has subsequently been used as the starting point for numerous AT schemes (Huang et al., 2015; Wong and Kolter, 2018; Kurakin et al., 2017; Madry et al., 2018; Goodfellow et al., 2015).
Pitfalls of surrogate-based optimization. Despite the intuitive appeal of this paradigm, surrogate-based adversarial attacks are known to overestimate robustness Mosbach et al. (2018); Croce et al. (2020b); Croce and Hein (2020), and standard adversarial training algorithms are known to fail against strong attacks. Furthermore, this formulation suffers from pitfalls such as robust overfitting (Rice et al., 2020) and trade-offs between robustness and accuracy (Tsipras et al., 2019b; Robey et al., 2022). To combat these shortcomings, empirical adversarial attacks and defenses have increasingly relied on heuristics such as multiple restarts, variable learning rate schedules (Croce and Hein, 2020), and carefully crafted initializations, resulting in a widening gap between the theory and practice of adversarial learning. In the next section, we argue that these pitfalls can be attributed to the fundamental limitations of (8).
@@ 3.1 Fundamental limitations of surrogates in adversarial learning
  • Limitation I: Weak attackers. In the adversarial evaluation problem of (7), the adversary maximizes an upper bound on the classification error. This means that any solution $\eta^{\star}$ to (7) is not guaranteed to increase the classification error in (5), resulting in weakened adversaries which are misaligned with the goal of finding adversarial examples that fool DNNs. Indeed, when the surrogate is an upper bound on the classification error, the only conclusion about the adversarial perturbation $\eta^{\star}$ obtained from (7) and its true objective (5) is:
@@
    Notably, the right hand side of (9) can be arbitrarily large while the left hand side can simultaneously be equal to zero, i.e., the problem in (7) can fail to produce an adversarial example, even at optimality. Thus, while it is known empirically that attacks based on (7) tend to overestimate robustness (Croce and Hein, 2020; Croce et al., 2020b; Gowal et al., 2019), we show that this shortcoming is evident a priori.
@@
Both of these limitations arise directly by virtue of rewriting (7) and (8) with the surrogate loss $\ell$. Therefore, to summarize, there is a distinct tension between the efficient, yet misaligned paradigm of surrogate-based adversarial training with the principled, yet intractable paradigm of minimax optimization on the classification error. In the remainder of this section, we resolve this tension by decoupling the optimization problems of the adversary and the training algorithm.
@@ 3.2 Decoupling adversarial attacks and defenses
While this problem still constitutes a zero-sum game, the role of the attacker (the constraint in (11)) and the role of the defender (the objective in (10)) are now decoupled. From this perspective, the tension engendered by introducing surrogate losses is laid bare: the attacker ought to maximize a lower bound of the classification error (c.f., Limitation I), whereas the defender ought to minimize an upper bound on the classification error (c.f., Limitation II). This implies that to preserve guarantees on optimality, the attacker and defender must optimize separate objectives. In what follows, we discuss these objectives for the attacker and defender in detail.
@@
Note that this does not constitute a relaxation; (11) and (14) are equivalent optimization problems. This means that the attacker can maximize the classification error directly using first-order optimization methods without resorting to a relaxation.
@@ 4 Algorithms
             $\eta_{j}\leftarrow\text{Unif}[\max(X-\epsilon,0),\min(X+\epsilon,1)]$ // (assume images in $[0,1]^{d}$)
@@
                   $\eta_{j}\leftarrow\text{OPTIM}(\eta_{j},\nabla_{\eta_{i}}M_{\theta}(x+\eta_{j},y)_{j})$ // (optimizer step, e.g., RMSprop)
                   $\eta_{j}\leftarrow\Pi_{B_{\epsilon}(X)\cap[0,1]^{d}}(\eta_{j})$ // (Projection onto valid perturbation set)
@@
             Sample $i\sim\text{Unif}[n]$ $\eta^{\star}\leftarrow\text{BETA}(x_{i},y_{i},\epsilon,f_{\theta},T^{\prime})$ $L(\theta)\leftarrow\ell(f_{\theta}(x_{i}+\eta^{\star}),y_{i})$ $\theta\leftarrow\text{OPTIM}(\theta,\nabla L(\theta))$ // (optimizer step, e.g., SGD)
@@
Given the non-zero-sum formulation of AT in the previous section, the next question is how one should solve this bilevel optimization problem in practice. Our starting point is the empirical version of this bilevel problem, wherein we assume access to a finite dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ of $n$ instance-label pairs sampled i.i.d. from $\mathcal{D}$.
@@
To solve this empirical problem, we adopt a stochastic optimization based approach. That is, we first iteratively sample mini-batches from our dataset uniformly at random, and then obtain adversarial perturbations by solving the lower level problems in (21) and (22). Note that given the differentiability of the negative margin, the lower level problems can be solved iteratively with generic optimizers, e.g., Adam Kingma and Ba (2014) or RMSprop. This procedure is summarized in Algorithm 1, which we call the BEst Targeted Attack (BETA), given that it directly maximizes the classification error.
After obtaining such perturbations, we calculate the perturbed loss in (20), and then differentiate through this loss with respect to the model parameters. By updating the model parameters $\theta$ in the negative direction of this gradient, our algorithm seeks classifiers that are robust against perturbations found by BETA. We call the full adversarial training procedure based on this attack BETA Adversarial Training (BETA-AT), as it invokes BETA as a subroutine; see Algorithm 2 for details.
Smoothing the lower level. One potential limitation of the BETA-AT algorithm introduced in Algorithm 2 is its sample efficiency: BETA computes one adversarial perturbation per class, but only one of these perturbations is chosen for the upper level of the bilevel formulation (20). In this way, one could argue that there is wasted computational effort in discarding perturbations that achieve high values of the negative margin (12). This potential shortcoming is a byproduct of the non-smoothness of the max operator in (22). Fortunately, we can alleviate this limitation by using smooth under-approximations of the max (e.g., the softmax function), which is continuously differentiable. We explore this scheme in Appendix B.

### v2

@@ Abstract
One prominent approach toward resolving the adversarial vulnerability of deep neural networks is the two-player zero-sum paradigm of adversarial training, in which predictors are trained against adversarially chosen perturbations of data. Despite the promise of this approach, algorithms based on this paradigm have not engendered sufficient levels of robustness and suffer from pathological behavior like robust overfitting. To understand this shortcoming, we first show that the commonly used surrogate-based relaxation used in adversarial training algorithms voids all guarantees on the robustness of trained classifiers. The identification of this pitfall informs a novel non-zero-sum bilevel formulation of adversarial training, wherein each player optimizes a different objective function. Our formulation yields a simple algorithmic framework that matches and in some cases outperforms state-of-the-art attacks, attains comparable levels of robustness to standard adversarial training algorithms, and does not suffer from robust overfitting.
@@ 1 Introduction
A longstanding disappointment in the machine learning (ML) community is that deep neural networks (DNNs) remain vulnerable to seemingly innocuous changes to their input data, including nuisances in visual data (Laidlaw et al., 2020; Hendrycks & Dietterich, 2019), sub-populations (Santurkar et al., 2021; Koh et al., 2021), and distribution shifts (Xiao et al., 2021; Arjovsky et al., 2019; Robey et al., 2021). Prominent amongst these vulnerabilities is the setting of adversarial examples, wherein it has been conclusively shown that imperceptible, adversarially-chosen perturbations can fool state-of-the-art classifiers parameterized by DNNs (Szegedy et al., 2013; Biggio et al., 2013). In response, a plethora of research has proposed so-called adversarial training (AT) algorithms (Madry et al., 2018; Goodfellow et al., 2015), which are designed to improve robustness against adversarial examples.
@@
AT is ubiquitously formulated as a two-player zero-sum game, where both players—often referred to as the defender and the adversary—respectively seek to minimize and maximize the classification error. However, this zero-sum game is not implementable in practice as the discontinuous nature of the classification error is not compatible with first-order optimization algorithms. To bridge this gap between theory and practice, it is commonplace to replace the classification error with a smooth surrogate loss (e.g., the cross-entropy loss) which is amenable to gradient-based optimization (Madry et al., 2018; Zhang et al., 2019). And while this seemingly harmless modification has a decades-long tradition in the ML literature due to the guarantees it imparts on non-adversarial objectives (Bartlett et al., 2006; Shalev-Shwartz & Ben-David, 2014; Roux, 2017), there is a pronounced gap in the literature regarding the implications of this relaxation on the standard formulation of AT.
@@
As the field of robust ML has matured, surrogate-based AT algorithms have collectively resulted in steady progress toward stronger attacks and robust defenses (Croce et al., 2020a). However, despite these advances, recent years have witnessed a plateau in robustness measures on popular leaderboards, resulting in the widely held beliefs that robustness and accuracy may be irreconcilable (Tsipras et al., 2019; Dobriban et al., 2020) and that robust generalization requires significantly more data (Schmidt et al., 2018; Chen et al., 2020). Moreover, various phenomena such as robust overfitting (Rice et al., 2020) have indicated that progress has been overestimated (Croce & Hein, 2020). To combat these pitfalls, state-of-the-art algorithms increasingly rely on ad-hoc regularization schemes (Kannan et al., 2018; Chan et al., 2020), weight perturbations (Wu et al., 2020; Sun et al., 2021), and heuristics such as multiple restarts, carefully crafted learning rate schedules, and convoluted stopping conditions, all of which contribute to an unclear set of best practices and a growing literature concerned with identifying flaws in various AT schemes (Latorre et al., 2023).
@@
Motivated by these challenges, we argue that the pervasive surrogate-based zero-sum approach to AT suffers from a fundamental flaw. Our analysis of the standard minimax formulation of AT reveals that maximizing a surrogate like the cross-entropy provides no guarantee that the the classification error will increase, resulting in weak adversaries and ineffective AT algorithms. In identifying this shortcoming, we prove that to preserve guarantees on the optimality of the classification error objective, the defender and the adversary must optimize different objectives, resulting in a non-zero-sum game. This leads to a novel, yet natural bilevel formulation (Bard, 2013) of AT in which the defender minimizes an upper bound on the classification error, while the attacker maximizes a continuous reformulation of the classification error. We then propose an algorithm based on our formulation which is free from heuristics and ad hoc optimization techniques. Our empirical evaluations reveal that our approach matches the test robustness achieved by the state-of-the-art, yet highly heuristic approaches such as AutoAttack, and that it eliminates robust overfitting.
Contributions. Our contributions are as follows.
@@
  • New adversarial training algorithm. We derive BETA, a new, heuristic-free algorithm based on our bilevel formulation which offers competitive empirical robustness on CIFAR-10.
  • Elimination of robust overfitting. Our algorithm does not suffer from robust overfitting. This suggest that robust overfitting is an artifact of the use of improper surrogates in the original AT paradigm, and that the use of a correct optimization formulation is enough to solve it.
  • State-of-the-art robustness evaluation. We show that our proposed optimization objective for the adversary yields a simple algorithm that matches the performance of the state-of-the-art, yet highly complex AutoAttack method, on state-of-the-art robust classifiers trained on CIFAR-10.
@@ 2.1 Preliminaries: Training DNNs with surrogate losses
Among the barriers to solving (1) in practice is the fact that the classification error is a discontinuous function of $\theta$, which in turn renders continuous first-order methods intractable. Fortunately, this pitfall can be resolved by minimizing a surrogate loss function $\ell:[k]\times[k]\to\mathbb{R}$ in place of the classification error (Shalev-Shwartz & Ben-David, 2014, §12.3). For minimization problems, surrogate losses are chosen to be differentiable upper bounds of the classification error of $f_{\theta}$ in the sense that
@@ 2.2 The pervasive setting of adversarial examples
For common hypothesis classes, it is well-known that classifiers obtained by solving (3) are sensitive to adversarial examples (Szegedy et al., 2013; Biggio et al., 2013), i.e., given an instance label pair $(X,Y)$, it is relatively straightforward to find perturbations $\eta\in\mathbb{R}^{d}$ with small norm $\left|\left|\eta\right|\right|\leq\epsilon$ for some fixed $\epsilon>0$ such that
@@ 2.3 Surrogate-based approaches to robustness
As discussed in § 2.1, the discontinuity of the classification error complicates the task of finding adversarial examples, as in (5), and of training against these perturbed instances, as in (6). One appealing approach toward overcoming this pitfall is to simply deploy a surrogate loss in place of the classification error inside (6), which gives rise to the following pair of optimization problems: $\eta^{\star}\in\operatorname*{arg\,max}_{\eta:\left|\left|\eta\right|\right|\leq\epsilon}\ell(f_{\theta}(X+\eta),Y)$ (7) $\min_{\theta\in\Theta}\mathbb{E}\left[\max_{\eta:\|\eta\|\leq\epsilon}\ell(f_{\theta}(X+\eta),Y)\right]$ (8)
Indeed, this surrogate-based approach is pervasive in practice. Madry et al.’s seminal paper on the subject of adversarial training employs this formulation (Madry et al., 2018), which has subsequently been used as the starting point for numerous AT schemes (Huang et al., 2015; Kurakin et al., 2017).
Pitfalls of surrogate-based optimization. Despite the intuitive appeal of this paradigm, surrogate-based adversarial attacks are known to overestimate robustness (Mosbach et al., 2018; Croce et al., 2020b; Croce & Hein, 2020), and standard adversarial training algorithms are known to fail against strong attacks. Furthermore, this formulation suffers from pitfalls such as robust overfitting (Rice et al., 2020) and trade-offs between robustness and accuracy (Zhang et al., 2019). To combat these shortcomings, empirical adversarial attacks and defenses have increasingly relied on heuristics such as multiple restarts, variable learning rate schedules (Croce & Hein, 2020), and carefully crafted initializations, resulting in a widening gap between the theory and practice of adversarial learning. In the next section, we argue that these pitfalls can be attributed to the fundamental limitations of (8).
@@ 3 Non-zero-sum formulation of adversarial training
Limitation I: Weak attackers. In the adversarial evaluation problem of (7), the adversary maximizes an upper bound on the classification error. This means that any solution $\eta^{\star}$ to (7) is not guaranteed to increase the classification error in (5), resulting in adversaries which are misaligned with the goal of finding adversarial examples. Indeed, when the surrogate is an upper bound on the classification error, the only conclusion about the perturbation $\eta^{\star}$ obtained from (7) and its true objective (5) is:
@@
Notably, the RHS of (9) can be arbitrarily large while the left hand side can simultaneously be equal to zero, i.e., the problem in (7) can fail to produce an adversarial example, even at optimality. Thus, while it is known empirically that attacks based on (7) tend to overestimate robustness (Croce & Hein, 2020), this argument shows that this shortcoming is evident a priori.
@@
Both of Limitation I and Limitation II arise directly by virtue of rewriting (7) and (8) with the surrogate loss $\ell$. To illustrate this more concretely, consider the following example.
Example 1.
Let $\epsilon>0$ be given, let $K$ denote the number of classes in a classification problem, and let $\ell$ denote the cross-entropy loss. Consider two possible logit vectors of class probabilities:
$z_{A}=(1/K+\epsilon,1/K-\epsilon,1/K,\ldots,1/K),\qquad z_{B}=(0.5-\epsilon,0.5+\epsilon,0,\ldots,0)$ (10)
Assume without loss of generality that the correct class is the first class. Then $z_{A}$ does not lead to an adversarial example, whereas $z_{B}$ does. However, observe that $\ell(z_{A},1)=-\log(1/K+\epsilon)$, which tends to $\infty$ as $K\to\infty$ and $\epsilon\to 0$. In contrast, $\ell(z_{B},1)=-\log(0.5-\epsilon)$ which remains bounded as $\epsilon\to 0$. Hence, an adversary maximizing the cross-entropy will always choose $z_{A}$ over $z_{B}$ and will therefore fail to identify the adversarial example.
Therefore, to summarize, there is a distinct tension between the efficient, yet misaligned paradigm of surrogate-based adversarial training with the principled, yet intractable paradigm of minimax optimization on the classification error. In the remainder of this section, we resolve this tension by decoupling the optimization problems of the attacker and the defender.
@@ 3.1 Decoupling adversarial attacks and defenses
While this problem still constitutes a zero-sum game, the role of the attacker (the constraint in (12)) and the role of the defender (the objective in (11)) are now decoupled. From this perspective, the tension engendered by introducing surrogate losses is laid bare: the attacker ought to maximize a lower bound on the classification error (c.f., Limitation I), whereas the defender ought to minimize an upper bound on the classification error (c.f., Limitation II). This implies that to preserve guarantees on optimality, the attacker and defender must optimize separate objectives. In what follows, we discuss these objectives for the attacker and defender in detail.
@@
Note that this does not constitute a relaxation; (12) and (15) are equivalent optimization problems. This means that the attacker can maximize the classification error directly using first-order optimization methods without resorting to a relaxation. Furthermore, in Appendix D, we give an example of a scenario wherein solving (15) retrieves the optimal adversarial perturbation whereas maximizing the standard adversarial surrogate fails to do so.
@@ 4 Algorithms
             $\eta_{j}\leftarrow\text{Unif}[\max(X-\epsilon,0),\min(X+\epsilon,1)]$ // Images are in $[0,1]^{d}$
@@
                   $\eta_{j}\leftarrow\text{OPTIM}(\eta_{j},\nabla_{\eta_{i}}M_{\theta}(x+\eta_{j},y)_{j})$ // Optimization step
                   $\eta_{j}\leftarrow\Pi_{B_{\epsilon}(X)\cap[0,1]^{d}}(\eta_{j})$ // Project onto perturbation set
@@
             Sample $i\sim\text{Unif}[n]$ $\eta^{\star}\leftarrow\text{BETA}(x_{i},y_{i},\epsilon,f_{\theta},T^{\prime})$ $L(\theta)\leftarrow\ell(f_{\theta}(x_{i}+\eta^{\star}),y_{i})$ $\theta\leftarrow\text{OPTIM}(\theta,\nabla L(\theta))$ // Optimization step
@@
Given the non-zero-sum formulation of AT, the next question is how one should solve this bilevel problem in practice. Our starting point is the empirical version of this bilevel problem, wherein we assume access to a finite dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ of $n$ instance-label pairs sampled i.i.d. from $\mathcal{D}$.
@@
To solve this empirical problem, we adopt a stochastic optimization based approach. That is, we first iteratively sample mini-batches from our dataset uniformly at random, and then obtain adversarial perturbations by solving the lower level problems in (22) and (23). Note that given the differentiability of the negative margin, the lower level problems can be solved iteratively with generic optimizers, e.g., Adam (Kingma & Ba, 2014) or RMSprop. This procedure is summarized in Algorithm 1, which we call the BEst Targeted Attack (BETA), given that it directly maximizes the classification error.
After obtaining such perturbations, we calculate the perturbed loss in (21), and then differentiate through this loss with respect to the model parameters. By updating the model parameters $\theta$ in the negative direction of this gradient, our algorithm seeks classifiers that are robust against perturbations found by BETA. We call the full adversarial training procedure based on this attack BETA Adversarial Training (BETA-AT), as it invokes BETA as a subroutine; see Algorithm 2 for details. Also see Figures LABEL:fig:timing,fig:performance in the appendix for an empirical study of the computational complexity of BETA.